# Quantum memory at nonzero temperature in a thermodynamically trivial system

### v1

@@ Abstract
Passive error correction protects logical information forever (in the thermodynamic limit) by updating the system based only on local information and few-body interactions. A paradigmatic example is the classical two-dimensional Ising model: a Metropolis-style Gibbs sampler retains the sign of the initial magnetization (a logical bit) for thermodynamically long times in the low-temperature phase. Known models of passive quantum error correction similarly exhibit thermodynamic phase transitions to a low-temperature phase wherein logical qubits are protected by thermally stable topological order. Here, in contrast, we show that constant-rate classical and quantum low-density parity check codes have no thermodynamic phase transitions at nonzero temperature, but nonetheless exhibit ergodicity-breaking dynamical transitions: below a critical nonzero temperature, the mixing time of local Gibbs sampling diverges in the thermodynamic limit. We conjecture that the circuit complexity of preparing extensive-energy states may diverge without crossing any thermodynamic transition. Fault-tolerant passive decoders, inspired by Gibbs samplers, may be amenable to measurement-free quantum error correction and may present a desirable experimental alternative to conventional quantum error correction based on syndrome measurements and active feedback.
@@ 1 Introduction
There is a deep relationship between error correction and thermodynamics. Data storage in magnetic devices (e.g. hard disk drives) relies on the fact that there are two stable magnetic configurations in a two-dimensional ferromagnet, such as the Ising model. Whether the spins point up or down determines whether a classical bit of information is 0 or 1. Crucially, the ferromagnetic phase is robust to thermal noise: at low temperatures, errors are passively corrected by the environment! The time until the stored data becomes corrupted is exponentially large in the system size Thomas (1989).
@@
The simplest analogy to the above story for a quantum error-correcting code is the four-dimensional toric code Dennis et al. (2002). Intuitively, it corresponds to “two copies of the 2D Ising model”: since a quantum code must protect against both $X$ (bit-flip) and $Z$ (phase-flip) errors, half of the dimensions of the 4D toric code protect against each. Since the 2D Ising model exhibits a phase transition to ferromagnetism below a nonzero critical temperature Onsager (1944), the 4D toric code also has a nonzero critical temperature below which the thermal state has topological order Alicki et al. (2008). This topological order can protect logical quantum information forever in the thermodynamic limit, even in the presence of thermal noise. In other words, the 4D toric code is a passive quantum memory.
@@
In all known examples of passive quantum error correction, there is a thermodynamic phase transition at nonzero temperature to a topologically ordered phase Alicki et al. (2008); Yoshida (2011); Hastings (2011); Brown et al. (2016); Liu and Lieu (2024). Such transitions are only known to exist in at least four spatial dimensions. Unfortunately, we live in three spatial dimensions. Therefore, many existing attempts to realize quantum error correction in experiment involve active decoding, where explicit measurements and feedback are necessary to protect quantum information. Yet many active decoders require global information about measurement outcomes, so the cost of performing such classical computations will increase as one builds ever-larger quantum computers. In contrast, a passively-decodable system corrects its own errors simply by thermalizing with a cold environment – far simpler, at least in principle, to realize experimentally.
@@
Motivated by the above challenges, this paper addresses the question of whether quantum memory can exist without any thermal phase transitions to a topologically ordered phase. An answer to this question is important, both because it could constrain the space of quantum codes to explore in which passive error correction is even feasible, and because it is important to understand whether the limited nonlocality Xu et al. (2023); Hong et al. (2023) that can be realized in quantum hardware is sufficient to realize passive quantum error correction.
@@
We will prove that a thermodynamic phase transition, at any nonzero temperature, is not necessary for finite-temperature self-correction in classical codes, and that a thermodynamic phase transition to a topologically ordered phase is not necessary for self-correction in quantum codes. Our analysis is of classical and quantum low-density parity-check (LDPC) codes Gallager (1962); Breuckmann and Eberhardt (2021). “Parity check” refers to the multi-bit generalizations of the ferromagnetic interactions in the Ising model; like in the Ising model, the interactions are frustration-free – they can all be satisfied simultaneously if all bits are in the 0 state (all spins up). Furthermore, the Hamiltonian of an LDPC code is $\mathsf{k}$-local – each term in the Hamiltonian only couples $\mathsf{k}=\mathrm{O}(1)$ degrees of freedom in the thermodynamic limit. However, there is no constraint on the physical range of these interactions when the interacting degrees of freedom are embedded in physical (three-dimensional) space. In fact, a typical LDPC code cannot be locally embedded in any finite dimension. The parity checks are “low-density” because they are $\mathsf{k}$-local, and each bit only participates in a finite number of parity checks. In this respect, these models are similarly “few-body” to the simpler Ising model, but can avoid stringent constraints on quantum error correction in few spatial dimensions Bravyi et al. (2010).
@@
The “infinite-dimensionality” of LDPC codes naturally leads to two properties, each of which are critical for our story. Firstly, LDPC codes can be non-redundant – every single parity check is an independent constraint from the others. This is not true in the two-dimensional Ising model, where the product of interactions around a square plaquette is always $+1$. This non-redundancy ensures thermodynamic triviality of the random LDPC codes that we study. Secondly, a random LDPC code typically has rapidly growing energy barriers to small perturbations, implying that the ground states of its corresponding Hamiltonian reside in very deep local minima in the energy landscape. The large energy barriers to escape these local minima ensure that Gibbs sampling Metropolis et al. (1953); Hastings (1970), a specific kind of passive error correction, is capable of trapping the system near its starting state, which enables the eventual decodability of the information.
@@ 2 Classical codes
The $2^{k}\geq 2^{n-m}$ ground states of the above Hamiltonian are called logical codewords, generated by $k$ logical bits. Notice that one codeword ($\mathcal{H}=-m$) is the state $Z_{i}=1$ for all $i$. We restrict our focus to codes in which each bit $Z_{i}$ is contained in exactly $\Delta_{B}$ checks, and take $\Delta_{B}<\Delta_{C}.$ It is conventional in the coding literature to organize the data above as follows: letting $\mathbb{F}_{2}=\{0,1\}$, we consider an $\mathbb{F}_{2}$-valued $m\times n$ matrix $H$ such that
@@
Defining the vector $\mathbf{z}\in\mathbb{F}_{2}^{n}$ with entries $z_{i}=\frac{1}{2}(1-Z_{i})$, we can write
@@
In other words, the number of logical bits scales linearly with the system size $n$. (2) There exist O(1) numbers $\gamma,\eta>0$ such that for all $\mathbf{z}$ obeying $0<|\mathbf{z}|\leq\gamma n$, $|H\mathbf{z}|\geq\eta|\mathbf{z}|$. This property is called “linear confinement”: a continuous path of bit flips between codewords will necessarily cross an extensive energy barrier. Since this confinement holds up to $\gamma n$, the code distance must also scale linearly with the system size.
Using these facts and following Yoshida (2011); Freedman and Hastings (2013); Weinstein et al. (2019), we now show that this LDPC code has a trivial free energy in the thermodynamic limit. Observe that since $k=n-m$, $H$ has no left null vectors. Hence, given any $\mathbf{s}\in\mathbb{F}_{2}^{m}$, there are exactly $2^{k}$ candidate states $\mathbf{z}$ satisfying $\mathbf{s}=H\mathbf{z}$. It is then straightforward to evaluate
@@
Let us now study the dynamics of Gibbs sampling for this LDPC code. We consider the discrete-time Metropolis algorithm in which
$\frac{\mathbf{P}[\mathbf{z}\rightarrow\mathbf{z}^{\prime}]}{\mathbf{P}[\mathbf{z}^{\prime}\rightarrow\mathbf{z}]}=\min\left(1,\mathrm{e}^{-\beta(\mathcal{H}(\mathbf{z}^{\prime})-\mathcal{H}(\mathbf{z}))}\right)\,.$ (7)
We study $\mathsf{k}$-local dynamics in which $|\mathbf{z}-\mathbf{z}^{\prime}|$ is O(1) when $\mathbf{P}[\mathbf{z}\rightarrow\mathbf{z}^{\prime}]\neq 0$. Suppose that at time $t=0$, we start in the codeword $\mathbf{0}$. To bound the time it takes for the Gibbs sampler to destroy information, it suffices to bound the time it takes for the Gibbs sampler to reach a state $\mathbf{z}_{t}\in E$, where the “bottleneck” set $E$ consists of all $|\mathbf{z}_{t}|=\gamma n/2<d/2$. In thermal equilibrium, we observe that
@@
How is it possible to have thermodynamic triviality, while the Gibbs sampler simultaneously protects a codeword? The answer lies in the peculiar energy landscape of these LDPC codes, illustrated in Figure 1. Locally near each codeword, the linear confinement property guarantees a very deep minimum in the energy landscape. If we could restrict $Z(\beta)$ to only count states $\mathbf{x}$ with $|\mathbf{x}|\leq\gamma n$, the Helmholtz free energy would exhibit a phase transition: above a critical temperature, the typical state would have $|\mathbf{x}|/n\rightarrow\gamma$, while below it the typical state is much closer to the codeword $\mathbf{0}$, as guaranteed by (8). In reality however, $Z(\beta)$ counts the whole state space, and when we look beyond the local landscape near a given codeword, we see a “swamp” of many very deep minima, some of which contain other codewords, but far more of which contain “sick” configurations where, e.g., a single parity check is flipped Rakovszky and Khemani (2023, 2024). Indeed, we have seen that there must exist a state with a single flipped parity-check since $H$ has no left null vectors; at the same time, such a state $\mathbf{x}$ obeys $|\mathbf{x}|=\mathrm{O}(n)$, and it too will have an extremely deep energy well where the Gibbs sampler would be stuck. The key point is the conceptual distinction between static equilibrium and dynamical equilibration – the thermodynamic free energy detects all of the exponentially many fake codewords, while the Gibbs sampler is stuck near whichever minimum it starts in.
To have a genuine thermodynamic phase transition, redundancy must be added to the parity checks, such that the sick configurations of Figure 1 gain finite energy density. The energetic suppression of these sick configurations is encapsulated by a closely related property to confinement known as soundness, which stipulates that small energy violations can always be produced by small errors. Note that the LDPC codes we have discussed are confined but not sound due to the low-energy sick configurations. The 2D Ising model, on the other hand, is both confined and sound. The intimate relationship between redundancy and soundness has been studied in the context of locally testable codes (LTCs) Ben-Sasson et al. (2009), in which one can reliably determine whether a state is a codeword by only querying a small number of physical bits. “Good” locally testable codes with constant rate, constant relative distance and constant query complexity ($c^{3}$-LTCs) have recently been found Panteleev and Kalachev (2022); Lin and Hsieh (2022); Dinur et al. (2022a). While redundancy and soundness is a feature of these codes, we have seen that they are unnecessary to have ergodicity-breaking at low temperatures.
@@ 3 Quantum codes
It is quite challenging to find a quantum LDPC code of $N$ qubits in which the code distance $D=\mathrm{O}(N)$, although recently some have been found Panteleev and Kalachev (2022); Leverrier and Zémor (2022); Dinur et al. (2022b). To keep the discussion as transparent as possible, we will instead discuss the simplest family of quantum LDPC codes, called hypergraph product (HGP) codes Tillich and Zemor (2014). HGP codes are formed by a graph product of two classical codes, each of which we take to be a good random LDPC code of the kind previously mentioned with parameters $[n,k=\mathrm{O}(n),d=\mathrm{O}(n)]$. The HGP construction is illustrated in Figure 2, while explicit formulas are relegated to Appendix B.1. One can prove the following important facts about such random HGP codes, each of which holds almost surely in the thermodynamic limit: (1) they are constant-rate codes with $N=n^{2}+(n-k)^{2}$, $D=d$, and $K/N=\mathfrak{r}^{2}$, with $\mathfrak{r}$ inherited from the classical LDPC code and given in (5). There are $M=n(n-k)$ checks of both $X$ and $Z$ type, and there are no redundant checks. (2) Like classical LDPC codes, they exhibit linear confinement Leverrier et al. (2015): given any binary string $\mathbf{x}\in\mathbb{F}^{N}_{2}$ corresponding to an $X$-type or $Z$-type error with weight $|\mathbf{x}|\leq\gamma n$, we have $|H_{X}\mathbf{x}|,|H_{Z}\mathbf{x}|\geq\eta|\mathbf{x}|$ for some $\eta=\mathrm{O}(1)$. (3) Every stabilizer involves exactly $\Delta_{B}+\Delta_{C}$ qubits, and each qubit is involved in at most $\Delta_{C}$ $X$-stabilizers and $\Delta_{C}$ $Z$-stabilizers.
@@
We understand the existence of this ergodicity-breaking in Gibbs sampling in exactly the same way as illustrated for classical LDPC codes. The quantum LDPC codes have a very complex energy landscape, such that the Gibbs sampler will get stuck in a very deep energy minimum for exponentially long times; in contrast, the thermodynamics is sensitive to the “swamp” of exponentially many deep local minima corresponding to very large errors that flip few parity checks. A thermal phase transition requires redundancy which our HGP codes do not have. Numerical simulations had previously suggested that passive decoding might be possible even above a nonzero critical temperature for topological order Hastings et al. (2014); we stress that in our construction, the model is completely thermodynamically trivial. Nonvanishing energy barriers have been the primary features of Haah’s cubic code Haah (2011) and welded codes Michnicki (2014). Nonetheless, it has been shown that the lifetime of logical information in these codes is finite at any positive temperature due to entropic effects Bravyi and Haah (2013); Siva and Yoshida (2017).
We have shown that passive error correction is possible for quantum LDPC codes. Previous work has shown that such codes have good greedy decoders Leverrier et al. (2015) and that single-shot quantum error correction Bombin (2015); Campbell (2019) – where a single round of noisy syndrome measurement suffices to suppress errors – is achievable Quintavalle et al. (2021); Fawzi et al. (2018); Gu et al. (2023). Passive error correction is arguably even more desirable than single-shot error correction, as the single-shot decoder may nevertheless require global information as part of the decoding, whereas by construction the Gibbs sampling dynamics described above does not. In other words, an autonomous machine can detect and correct quantum errors in a memoryless way. Below a physical error threshold, the logical error rate can be suppressed to arbitrarily low levels by increasing the size of the system. We emphasize that since the quantum LDPC codes have a nonvanishing rate in the thermodynamic limit, the overhead (number of physical qubits per logical qubit) remains constant.

### v2

@@ Abstract
Passive error correction protects logical information forever (in the thermodynamic limit) by updating the system based only on local information and few-body interactions. A paradigmatic example is the classical two-dimensional Ising model: a Metropolis-style Gibbs sampler retains the sign of the initial magnetization (a logical bit) for thermodynamically long times in the low-temperature phase. Known models of passive quantum error correction similarly exhibit thermodynamic phase transitions to a low-temperature phase wherein logical qubits are protected by thermally stable topological order. Here, in contrast, we show that certain families of constant-rate classical and quantum low-density parity check codes have no thermodynamic phase transitions at nonzero temperature, but nonetheless exhibit ergodicity-breaking dynamical transitions: below a critical nonzero temperature, the mixing time of local Gibbs sampling diverges in the thermodynamic limit. Slow Gibbs sampling of such codes enables fault-tolerant passive quantum error correction using finite-depth circuits. This strategy is well suited to measurement-free quantum error correction, and may present a desirable experimental alternative to conventional quantum error correction based on syndrome measurements and active feedback.
@@ 1 Introduction
There is a deep relationship between error correction and thermodynamics. Data storage in magnetic devices (e.g. hard disk drives) relies on the fact that there are two stable magnetic configurations in a ferromagnet, such as the Ising model, in $d\geq 2$ spatial dimensions. Whether the spins point up or down determines whether a classical bit of information is 0 or 1. Crucially, the ferromagnetic phase is robust to thermal noise: at low temperatures, errors are passively corrected by the environment itself! The time until the stored data becomes corrupted is exponentially long in the system size Thomas (1989).
@@
The simplest analogy to the above story for a quantum error-correcting code is the four-dimensional toric code Dennis et al. (2002). Intuitively, it corresponds to “two copies of the 2D Ising model”: since a quantum code must protect against both $X$ (bit-flip) and $Z$ (phase-flip) errors, half of the dimensions of the 4D toric code protect against each. Similar to how the 2D Ising model exhibits ferromagnetic order below a nonzero critical temperature Onsager (1944), the 4D toric code also has a nonzero critical temperature below which its thermal state contains topological order Alicki et al. (2008). This topological order can protect logical quantum information forever in the thermodynamic limit, even in the presence of thermal noise. In other words, the 4D toric code is a passive quantum memory.
@@
In all known examples of passive quantum error correction, there is a thermodynamic phase transition at nonzero temperature to a topologically ordered phase Alicki et al. (2008); Yoshida (2011); Hastings (2011); Brown et al. (2016); Liu and Lieu (2024). Such transitions are only known to exist in at least four spatial dimensions. Unfortunately, we live in three spatial dimensions. Therefore, many existing attempts to realize quantum error correction in experiment involve active decoding, where projective measurements and feedback are necessary to protect quantum information. These active decoders typically require global information about measurement outcomes, and so the cost of performing such classical computations will increase as one builds ever-larger quantum computers. In contrast, a passively-decodable system corrects its own errors simply by thermalizing with a cold environment – far simpler, at least in principle, to implement experimentally.
@@
Motivated by the above challenges, we address the question of whether quantum memory can exist without any thermodynamic transition to a topologically ordered phase. An answer to this question is important, both because it will constrain the landscape of passively-correctable quantum codes, and because it is important to understand whether the limited nonlocality Xu et al. (2023); Hong et al. (2023) that can be realized in quantum hardware will enable passive quantum error correction.
@@
We will prove that a thermodynamic phase transition, at any nonzero temperature, is not necessary for finite-temperature self-correction in classical codes, and that a thermodynamic phase transition to a topologically ordered phase is not necessary for self-correction in quantum codes. Our analysis is of classical and quantum low-density parity-check (LDPC) codes Gallager (1962); Breuckmann and Eberhardt (2021). “Parity check” refers to the multi-bit generalizations of the ferromagnetic interactions in the Ising model; like in the Ising model, the interactions are frustration-free – they can all be satisfied simultaneously if all bits are in the 0 state (all spins up). Furthermore, the Hamiltonian of an LDPC code is $\mathsf{k}$-local – each term in the Hamiltonian only couples $\mathsf{k}=\mathrm{O}(1)$ degrees of freedom in the thermodynamic limit. However, there is no constraint on the physical range of these interactions when the interacting degrees of freedom are embedded in physical (three-dimensional) space. In fact, a typical LDPC code cannot be locally embedded in any finite dimension. The parity checks are “low-density” because they are $\mathsf{k}$-local, and each bit only participates in a finite number of parity checks. In this respect, these models are similarly “few-body” to the simpler Ising model, but can avoid stringent constraints on quantum error correction in low spatial dimensions Bravyi et al. (2010).
@@
The “infinite-dimensionality” of LDPC codes naturally leads to two properties, each of which are critical for our story. Firstly, LDPC codes can be non-redundant – every single parity check is an independent constraint from the others. This is not true in the two-dimensional Ising model, where the product of interactions around a square plaquette is always $+1$. This non-redundancy ensures thermodynamic triviality of the random LDPC codes that we study. Secondly, a LDPC codes can possess rapidly growing energy barriers to small perturbations Sipser and Spielman (1996), implying that the ground states of its corresponding Hamiltonian reside in very deep local minima in the energy landscape. The large energy barriers to escape these local minima ensure that Gibbs sampling Metropolis et al. (1953); Hastings (1970), a specific kind of passive error correction, is capable of trapping the system near its starting state, which enables the eventual decodability of the information.
Previous works on random, classical LDPC code ensembles have demonstrated both the thermodynamic triviality Montanari and Semerjian (2006a) as well as the existence of a dynamical transition reminiscent to those in spin glasses Kirkpatrick and Thirumalai (1987); Montanari and Semerjian (2006b). For completeness and pedagogy, we will derive these results explicitly. We can then more directly explain the main result of this paper, which is the separation between quantum self-correction and thermodynamic phase transitions, which is qualitatively similar (but quantitatively different) to the classical story.
@@ 2 Classical codes
Here $S_{j}$ denotes the $j^{\mathrm{th}}$ bit in $S$. The $2^{k}\geq 2^{n-m}$ ground states of the above Hamiltonian are called logical codewords, generated by $k$ logical bits. Notice that one codeword ($\mathcal{H}=-m$) is the state $Z_{i}=1$ for all $i$. We restrict our focus to codes in which each bit $Z_{i}$ is contained in exactly $\Delta_{B}$ checks, and take $\Delta_{B}<\Delta_{C}.$ It is conventional in the coding literature to organize the data above as follows: letting $\mathbb{F}_{2}=\{0,1\}$, we consider an $\mathbb{F}_{2}$-valued $m\times n$ matrix $H$ such that
@@
Defining vector $\mathbf{z}\in\mathbb{F}_{2}^{n}$ as $z_{i}=\frac{1}{2}(1-Z_{i})$,
@@
In other words, the number of logical bits scales linearly with the system size $n$. (2) There exist O(1) numbers $\gamma,\eta>0$ such that for all $\mathbf{z}$ obeying $0<|\mathbf{z}|\leq\gamma n$, $|H\mathbf{z}|\geq\eta|\mathbf{z}|$. This property is known as linear confinement, and it implies that a continuous path of bit flips between codewords will necessarily cross an extensive energy barrier. Since this confinement holds up to $\gamma n$, the code distance must also scale linearly with the system size.
Using these facts and following Montanari and Semerjian (2006a); Yoshida (2011); Freedman and Hastings (2013); Weinstein et al. (2019), we now show that this LDPC code has a trivial free energy in the thermodynamic limit. Observe that since $k=n-m$, $H$ has no left null vectors. Hence, given any $\mathbf{s}\in\mathbb{F}_{2}^{m}$, there are exactly $2^{k}$ candidate states $\mathbf{z}$ satisfying $\mathbf{s}=H\mathbf{z}$. It is then straightforward to evaluate
@@
Let us now study the dynamics of Gibbs sampling for this LDPC code. We consider the discrete-time Metropolis algorithm which flips one bit at a time:
$\mathbf{P}[\mathbf{z}\rightarrow\mathbf{z}^{\prime}\neq\mathbf{z}]=\frac{1}{n}\min\left(1,\mathrm{e}^{-\beta(\mathcal{H}(\mathbf{z}^{\prime})-\mathcal{H}(\mathbf{z}))}\right)\,\mathbb{I}\left(|\mathbf{z}-\mathbf{z}^{\prime}|=1\right)\,,$ (7)
where $\mathbb{I}(\cdot)$ is the indicator function which evaluates to 1 if its argument is true and 0 otherwise. Suppose that at time $t=0$, we start in the codeword $\mathbf{0}$. To bound the time it takes for the Gibbs sampler to destroy information, it suffices to bound the time it takes for the Gibbs sampler to reach a state $\mathbf{z}_{t}\in E$, where the “bottleneck” set $E$ consists of all $|\mathbf{z}_{t}|=\gamma n/2<d/2$. In thermal equilibrium, we observe that
@@
How is it possible to have thermodynamic triviality, while the Gibbs sampler simultaneously protects a codeword? The answer lies in the peculiar energy landscape of these LDPC codes, illustrated in Figure 1. Locally near each codeword, the linear confinement property guarantees a very deep minimum in the energy landscape. If we could restrict $Z(\beta)$ to only count states $\mathbf{x}$ with $|\mathbf{x}|\leq\gamma n$, the Helmholtz free energy would exhibit a phase transition: above a critical temperature, the typical state would have $|\mathbf{x}|/n\rightarrow\gamma$, while below it the typical state is much closer to the codeword $\mathbf{0}$, as guaranteed by (8). In reality however, $Z(\beta)$ counts the whole state space, and when we look beyond the local landscape near a given codeword, we see a “swamp” of many very deep minima, some of which contain other codewords, but far more of which contain “sick” configurations where, e.g., a single parity check is flipped Mézard and Montanari (2009); Rakovszky and Khemani (2023, 2024). Indeed, we have seen that there must exist a state with a single flipped parity-check since $H$ has no left null vectors; at the same time, such a state will be far from all codewords, and it too will have an extremely deep energy well where the Gibbs sampler would be stuck. The key point is the conceptual distinction between static equilibrium and dynamical equilibration – the thermodynamic free energy detects all of the exponentially many fake codewords, while the Gibbs sampler is stuck near whichever minimum it starts in.
To have a genuine thermodynamic phase transition, redundancy must be added to the parity checks, such that the sick configurations of Figure 1 gain finite energy density. The energetic suppression of these sick configurations is encapsulated by a closely related property to confinement known as soundness, which stipulates that small energy violations can always be produced by small errors. Note that the LDPC codes we have discussed are confined but not sound due to the low-energy sick configurations. The 2D Ising model, on the other hand, is both (sublinearly) confined and sound. The intimate relationship between redundancy and soundness has been studied in the context of locally testable codes (LTCs) Ben-Sasson et al. (2009), in which one can reliably determine whether a state is a codeword by only querying a small number of physical bits. “Good” locally testable codes with constant rate, constant relative distance and constant query complexity ($c^{3}$-LTCs) have recently been found Panteleev and Kalachev (2022); Lin and Hsieh (2022); Dinur et al. (2022a). While redundancy and soundness is a feature of these codes, we have seen that they are unnecessary to have ergodicity-breaking at low temperatures.
@@ 3 Quantum codes
It is quite challenging to find a quantum LDPC code of $N$ qubits in which the code distance $D=\mathrm{O}(N)$, although recently some have been found Panteleev and Kalachev (2022); Leverrier and Zémor (2022); Dinur et al. (2022b). To keep the discussion as transparent as possible, we will instead discuss the simplest family of quantum LDPC codes, called hypergraph product (HGP) codes Tillich and Zemor (2014). HGP codes are formed by a graph product of two classical codes, each of which we take to be a good random LDPC code of the kind previously mentioned with parameters $[n,k=\mathrm{O}(n),d=\mathrm{O}(n)]$. The HGP construction is illustrated in Figure 2, while explicit formulas are relegated to Appendix B.1. One can prove the following important facts about such random HGP codes, each of which holds almost surely in the thermodynamic limit: (1) they are constant-rate codes with $N=n^{2}+(n-k)^{2}$, $D=d$, and $K/N=\mathfrak{r}^{2}$, with $\mathfrak{r}$ inherited from the classical LDPC code and given in (5). There are $M=n(n-k)$ checks of both $X$ and $Z$ type, and there are no redundant checks. (2) Like classical LDPC codes, they can exhibit linear confinement Leverrier et al. (2015): given any binary string $\mathbf{x}\in\mathbb{F}^{N}_{2}$ corresponding to an $X$-type or $Z$-type error with weight $|\mathbf{x}|\leq\gamma n$, we have $|H_{X}\mathbf{x}|,|H_{Z}\mathbf{x}|\geq\eta|\mathbf{x}|$ for some $\eta=\mathrm{O}(1)$. (3) Every stabilizer involves exactly $\Delta_{B}+\Delta_{C}$ qubits, and each qubit is involved in at most $\Delta_{C}$ $X$-stabilizers and $\Delta_{C}$ $Z$-stabilizers.
@@
We understand the existence of this ergodicity-breaking in Gibbs sampling in exactly the same way as illustrated for classical LDPC codes. The quantum LDPC codes have a very complex energy landscape, such that the Gibbs sampler will get stuck in a very deep energy minimum for exponentially long times; in contrast, the thermodynamics is sensitive to the “swamp” of exponentially many deep local minima corresponding to very large errors that flip few parity checks. A thermal phase transition requires redundancy which our HGP codes do not have. Numerical simulations had previously suggested that passive decoding might be possible even above a nonzero critical temperature for topological order Hastings et al. (2014); we stress that in our construction, the model is completely thermodynamically trivial. Nonvanishing energy barriers have been the primary features of Haah’s cubic code Haah (2011) and welded codes Michnicki (2014). Nonetheless, the lifetime of logical information in these codes is finite at any positive temperature due to entropic effects Bravyi and Haah (2013); Siva and Yoshida (2017).
Despite the qualitative similarity between the coexistence of thermodynamic triviality, and low-temperature self-correction, classical self-correction likely does not directly imply self-correction in a quantum code formed from the hypergraph product of the classical codes: we provide both analytical and numerical arguments in Appendix D that there exist classical self-correcting codes whose hypergraph product is not a quantum self-correcting memory. This is because, as we explain, confinement in classical codes does not directly correspond to confinement in the hypergraph product quantum code.